Determine the basis for the set of vectors lying on the plane π‘Š = {(π‘₯, 𝑦, 𝑧): 3π‘₯ + 𝑦 βˆ’ 5𝑧 = 0}


1
Expert's answer
2022-03-03T17:47:50-0500

"y = -3x + 5z," then the points that are satisfying the equation of the plane:


"(x, -3x + 5z, z)."

We need to find two different vectors satisfying above that are not a multiple of each other.


"(x, -3x + 5z, z) = x(1, -3, 0) + z(0, 5, 1)."

An obvious basis:


"\begin{Bmatrix}
 (1, -3, 0), (0, 5, 1) 
\end{Bmatrix}."